# Specific heat of Gd3+- and Eu2+-based magnetic compounds

### v2

@@ VI Fits of experimental data
Since all the compounds mentioned above contain transition-metal ions, which in some cases are magnetic, one might question whether they influence the reported specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is very near ln(8) per rare-earth ion (consistent with $S=7/2$). This indicates that any effect from the transition metals is negligible.
Figure 6: (Color online) Specific heat over temperature as a function of temperature for several compounds and our fit. $K=0$ for the Gd compounds. Experimental data were taken from Refs. 29, 30, 32, 31, 33, 35, 34.
@@ VII Summary
Using quantum Monte Carlo, we have calculated the specific heat of a large number of non-frustrated magnetic structures. As expected from mean-field theory, the critical temperature $T_{C}$ increases approximately proportionally to $I=\Sigma_{\delta}z_{\delta}|J_{\delta}|$, where $J_{\delta}$ is the exchange interaction at distance $\delta$ and $z_{\delta}$ is the number of neighbors at this distance. However, $T_{C}$ lies below the mean-field value.
Rather surprisingly, we find that the specific heat depends mainly on the effective number of neighbors $z$ [see Eq. (3)] and is not quite sensitive to the particular structure. This allows us to derive general conclusions on the effect of $z$, independently of the particular structure. For large $z$, quantum fluctuations are reduced, and the specific heat approaches the mean-field value. Reducing $z$, the specific heat decreases below $T_{C}$ and increases above $T_{C}$ displaying larger tails.
For a quantitative comparison with experimental results, it is essential to incorporate axial anisotropy $H_{A}$ (see Section IV). The origin of $H_{A}$ lies in a small admixture of ${}^{6}P_{7/2}$ in the ground-state multiplet. This admixture introduces a component with $L=1$, rendering the system sensitive to axial crystal fields.
Using $z$ as the only free parameter in addition to the critical temperature, we have fitted the specific heat of four Gd compounds (GdNi3Ga9, GdPdBi, GdCu2Ge2, and GdNiSi3). The specific heat of GdCoIn5 has been fitted previously using the same approach [19]. Curiously, we find that $H_{A}$ can be neglected for these Gd compounds. Including also $K$ as a parameter, we have done the same fitting for two Eu compounds (Eu2Pd3Sn3 and EuPdSn2). The agreement shown in Fig. 6 is noticeable. It is possible that the larger atomic volume of Eu+2 renders this ion more susceptible to crystal-field effects.
Note that while most of our calculations have been done on non-frustrated systems, some of the fitted compounds here (GdPdBi and GdNiSi3), and GdCoIn5 fitted previously [19], are expected to have some degree of frustration.
Our work opens the possibility to classify and analyze the specific heat of non-frustrated Gd+3 and Eu+2, at least in a first semiquantitative manner using the above-discussed two parameters.
@@ Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. A.A.A. acknowledges financial support provided by PICT 2020A 03661 of the Agencia I+D+i, Argentina. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina.
@@ VIII Appendix
Here we calculate the mean value and fluctuations of a random variable $x$ with different intensities  [^1] The same result can be obtained considering each bond as an independent variable.. For simplicity we consider only two intensities $J_{1}$ and $J_{2}$. Extension to the general case is straightforward. We consider $z_{1}$ attempts in which the probability of the result $J_{1}$ is $p$ and that of $-J_{1}$ is $q=1-p$. In addition, there are $z_{2}$ attempts with probability $r$ of finding the result $J_{2}$ and $s=1-r$ for the result $-J_{2}$. At the end we will take $r=p$. For non-frustrated systems, both $J_{\delta}$ can be chosen positive. Clearly, the expectation value of $x$ is
@@
×$\times\left[nJ_{1}-(z_{1}-n)J_{1}+mJ_{2}-(z_{2}-m)J_{2}\right]^{2}.$
@@
Using Eqs. (9) and (13), setting $r=p$ and replacing $q=s=1-p$, one finally obtains

### v3

@@ VI Fits of experimental data
Since all the compounds contain transition-metal ions, which may be magnetic, one might question their influence on the specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is close to $\ln(8)$ per rare-earth ion (consistent with $S=7/2$), indicating that any effects from transition metals are negligible.
The success of these fits suggests that the effective number of neighbors $z$ is the dominant parameter determining the specific heat of 4f7 systems, with axial anisotropy $K$ providing additional refinement for Eu compounds. This approach significantly outperforms mean-field theory, which would fail to reproduce the observed specific heat curves (see Figure 2).
Figure 6: Experimental specific heat $C/T$ versus $T/T_{N}$ for six 4f7 compounds (dots) and our theoretical fits (lines). $K=0$ for all Gd-based compounds (GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3). Experimental data sources: Refs. 33, 34, 36, 35, 27, 29, 38.
@@ VII Summary
We present a unified framework for understanding the specific heat of 4f7 magnetic compounds (Gd3+ and Eu2+), characterized by a ground state with $J=7/2$ and $L=0$. Through extensive quantum Monte Carlo simulations, we demonstrate that the specific heat is primarily governed by two parameters: the effective number of neighbors $z$ and axial anisotropy $K$.
The effective number of neighbors $z$ quantifies the strength of spatial and quantum fluctuations, with larger $z$ values corresponding to reduced fluctuations and specific heat curves closer to mean-field predictions. Crucially, we find that the specific heat is remarkably insensitive to the detailed lattice structure for a given $z$, enabling a universal description of diverse magnetic systems.
The origin of the anisotropy $K$ lies in a small admixture of $^{6}P_{7/2}$ in the ground-state multiplet. This admixture introduces a component with $L=1$, rendering the system sensitive to axial crystal fields.
We successfully fit the specific heat of six experimental compounds (four Gd-based and two Eu-based) using only $z$ and $K$ as free parameters, with excellent agreement between theory and experiment. Notably, axial anisotropy is essential for Eu compounds but can be neglected for Gd compounds, reflecting the larger ionic radius of Eu2+ and its greater susceptibility to crystal-field effects. This work provides a simple yet powerful tool for interpreting specific heat data in 4f7 systems, significantly outperforming mean-field approaches and enabling the extraction of (effective) fundamental magnetic interaction parameters from experimental measurements. While our two-parameter model works well for the compounds studied, it may require additional parameters for systems with strong frustration or complex magnetic order.
Experimental researchers studying new 4f7 compounds can now follow this straightforward procedure: (1) Measure the specific heat and determine $T_{N}$; (2) Compare the $C/T$ vs $T/T_{N}$ curve to our reference data in Ref. [31]; (3) Identify the $z$ value that provides the best match; (4) Use Equation 4 and known crystal structure to constrain possible exchange interaction patterns. This approach provides physically meaningful constraints on magnetic interactions. These results could also guide the development of more accurate theoretical models for compounds.
@@ Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. A.A.A. acknowledges financial support provided by PICT 2020A 03661 of the Agencia I+D+i, Argentina. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina. Computational resources were provided by the HPC cluster of the Physics Department at Centro Atómico Bariloche (CNEA), which is part of Argentina’s National High-Performance Computing System (SNCAD).
@@ VIII Appendix
Here we calculate the mean value and fluctuations of a random variable $x$ with different intensities  [^1] The same result can be obtained considering each bond as an independent variable.. For simplicity we consider only two intensities $J_{1}$ and $J_{2}$. Extension to the general case is straightforward. We consider $z_{1}$ attempts in which the probability of the result $J_{1}$ is $p$ and that of $J_{1}$ is q=1pq=1-p. In addition, there are $z_{2}$ attempts with probability $r$ of finding the result $J_{2}$ and s=1rs=1-r for the result $J_{2}$. At the end we will take $r=p$. For non-frustrated systems, both $J_{\delta}$ can be chosen positive. Clearly, the expectation value of $x$ is
@@
$\times\left[nJ_{1}-(z_{1}-n)J_{1}+mJ_{2}-(z_{2}-m)J_{2}\right]^{2}.$
@@
Using Eqs. (10) and (14), setting $r=p$ and replacing q=s=1pq=s=1-p, one finally obtains